OpenStudy (anonymous):

15% of what number is 24?

OpenStudy (kropot72):

Let the number to be found be x. One percent of x = 24/15 One hundred percent of x = (24/15) x 100 = 160 x = 160

OpenStudy (anonymous):

\[24\div15\times100=160\]